OpenStudy (anonymous):

an airplane flew for 6 hours against a tail wind of 25 km/hr. the return trip with the same wind took 5 hours. find the speen of the airplane in still air

OpenStudy (lacypennelll):

5(p+25)=6(p-25) 5p+125=6p-150 p=275 km/hr Plane speed in still air = 275km/hr

OpenStudy (lacypennelll):

hope this helped